OpenStudy (anonymous):

cotθ= (square root of 3) 0ver 3

OpenStudy (john_es):

\[\cot\theta=\sqrt{3}/3\Rightarrow\tan\theta=3/\sqrt{3}\Rightarrow\tan\theta=\sqrt{3}\Rightarrow \theta=\arctan\sqrt{3}=60^0+360k; \ k=1,2,...\]

OpenStudy (john_es):

I mean k=1,2,...

OpenStudy (anonymous):

okay thank you so much!